A motorist starts from city A at 2:00 PM. And travels to city B at an average speed of 30mi/hr. After resting at B for one hour, she returns over the same route at an average speed of 40mi/hr and arrives at A that evening at 6:30 PM. Determine the distance between A and B. Please, Thank you~ <3
Answer by mananth(16949)

Let the distance be x miles
resting time = 1 hour
Total driving time = 4.5 -1 = 3.5 hours
Time onward + time return = 3.5
let d= x
t=d/r
x/30 +x/40 =3.5
multiply equation by 1200 (30 * 40)
40x +30x= 4200
70x = 4200
/70
x= 60 miles
HAPPY NEW YEAR